OpenStudy (anonymous):

7.00×10−3 mol of HBrare dissolved in water to make 12.0 L of solution. What is the concentration of hydroxide ions, [OH-], in this solution?

OpenStudy (matt101):

HBr dissociates into H+ and Br-. Therefore, if you have 7.00×10^(−3) mol of HBr, you have 7.00×10(-3) mol of H+. The concentration of H+ is: \[[H+]=\frac{7\times10^{-3}}{12} = 5.83\times10^{-4} M\] The pH of the solution is: \[pH = -\log({5.83\times10^{-4})}=3.23\] This means the pOH of the solution is 14 - 3.23 = 10.77. Now work backwards. Use the pOH to find the concentration of OH- ions: \[[OH^-]=10^{-10.77}=1.70\times10^{-11}M\]

OpenStudy (anonymous):

Thank You!